# Geological Everything Model 3D: A Promptable Foundation Model for Unified and Zero-Shot Subsurface Understanding

### v2

@@ Abstract
Subsurface understanding is critical for energy transition, natural hazard mitigation, and planetary science. Yet subsurface analysis remains fragmented, with separate models required for structural interpretation, stratigraphic analysis, geobody segmentation, and property modeling—each tightly coupled to specific data distributions and task formulations. We introduce the Geological Everything Model 3D (GEM), a unified generative architecture that reformulates all these tasks as prompt-conditioned inference along latent structural frameworks derived from subsurface imaging. This formulation moves beyond task-specific models by enabling a shared inference mechanism, where GEM propagates human-provided prompts—such as well logs, masks, or structural sketches—along inferred structural frameworks to produce geologically coherent outputs. Through this mechanism, GEM achieves zero-shot generalization across tasks with heterogeneous prompt types, without retraining for new tasks or data sources. This capability emerges from a two-stage training process that combines self-supervised representation learning on large-scale field seismic data with adversarial fine-tuning using mixed prompts and labels across diverse subsurface tasks. GEM demonstrates broad applicability across surveys and tasks, including Martian radar stratigraphy analysis, structural interpretation in subduction zones, full seismic stratigraphic interpretation, geobody segmentation, and property modeling. By bridging expert knowledge with generative reasoning in a structurally aware manner, GEM lays the foundation for scalable, human-in-the-loop geophysical AI—transitioning from fragmented pipelines to a vertically integrated, promptable reasoning system.
@@ 1 Introduction
Subsurface understanding is foundational to some of society’s most pressing challenges: sustainable energy exploration [1, 2], geological carbon storage [3, 4], natural hazard prediction [5, 6] and stratigraphy [7, 8]. In recent years, Artificial Intelligence (AI) has emerged as a transformative paradigm for subsurface analysis [9, 10, 11, 12, 13]. However, current implementations remain fragmented, brittle, and narrowly scoped. Most workflows depend on task-specific models—each trained on narrow datasets and bound to a single survey region—to perform structural interpretation [14, 15, 16], stratigraphic analysis[17, 18, 19], geobody segmentation [20, 21, 22, 23] or property modeling [24, 25, 26, 27]. This fragmentation hinders generalization and requires costly retraining for every new task or area. While foundation models have revolutionized vision and language[28, 29, 30, 31, 32], no such model has emerged for subsurface imaging. This is not simply a matter of scale, but of epistemology: geological imaging data is sparse, indirect, and structurally entangled—lacking the high-density semantic cues that foundation models in other fields rely on. Bridging this domain gap calls for rethinking how we represent, prompt, and reason about the planetary interior—not as a patchwork of disconnected labels, but as a unified structure-conditioned generative process.
@@
In response to these limitations, we introduce the Geological Everything Model 3D (GEM), a unified generative framework that reformulates subsurface interpretation as prompt-conditioned inference over latent structural representations. Instead of treating tasks like structural interpretation, geobody segmentation, or property modeling as separate problems requiring bespoke models, GEM abstracts them as different forms of dimensional completion guided by human-provided prompts—such as well logs, masks, or structural sketches. At its core, GEM learns to propagate these prompts through a structurally coherent latent space derived from the input subsurface volume, enabling geologically consistent outputs across diverse task types. This formulation unifies a wide range of interpretation goals within a single model, while maintaining flexibility for expert interaction and scalability across surveys. By shifting the modeling paradigm from task-specific classification to structure-aware generation, GEM offers a foundation for zero-shot, prompt-driven, human-in-the-loop geological reasoning.
@@
GEM is trained in two stages to enable broad generalization without task-specific retraining. First, a self-supervised pretraining stage leverages masked representation learning over more than 500 field seismic volumes to learn structure-sensitive features without requiring labels. Then, adversarial fine-tuning is performed using synthetic data with diverse prompt-label pairs across structural, stratigraphic, geobody, and physical property domains. This combination allows GEM to align prompts with latent geological frameworks, supporting flexible input types—including lines, masks, and well logs—and producing task-consistent outputs with minimal supervision. Once trained, GEM generalizes across surveys, data modalities, and interpretation goals, achieving zero-shot performance without architectural modification or additional fine-tuning. We evaluate GEM across a diverse set of tasks and geological settings, demonstrating broad applicability and strong zero-shot generalization. Without retraining, GEM performs high-resolution Martian radar stratigraphy analysis, interprets complex 3D fault systems in subduction zones, conducts full seismic stratigraphic interpretation, segments geobodies such as salt domes and channels, and models properties across surveys with sparse well control. These applications span planetary exploration, structural geology and property modeling—tasks that traditionally require disconnected models, domain-specific workflows, and extensive retraining at each stage. In contrast, GEM consolidates structural interpretation, stratigraphic analysis, geological modeling, and property prediction within a single generative framework. This unified approach eliminates the need for task-specific pipelines and mitigates error accumulation across stages, offering a more stable, scalable, and interpretable alternative. By integrating expert prompts with structure-aware reasoning, GEM bridges geoscientific insight and generative modeling, establishing a new paradigm for human-guided subsurface AI. This work makes three key contributions:
@@
  • First, it proposes a unified generative mechanism for subsurface interpretation by reframing structural delineation, full stratigraphic analysis, geobody segmentation, and property modeling as prompt-guided completions over latent geological frameworks derived from imaging data. This formulation enables structure-aware reasoning across tasks within a shared representation space.
@@
  • Third, it demonstrates GEM ’s strong zero-shot generalization and broad applicability across surveys, modalities, and interpretation goals, including Martian radar stratigraphy, 3D fault interpretation in subduction zones, full seismic stratigraphic interpretation, geobody segmentation, and property modeling with sparse well control—often superior to task-specific supervised models.
@@ 2.1 GEM: A Unified and Prompt-Driven Reasoning Framework for Subsurface imaging
GEM is designed as a unified, prompt-driven foundation model for diverse subsurface imaging tasks. It reformulates structural interpretation, stratigraphic analysis, geobody segmentation, and physical property modeling as conditional generative processes, driven by sparse user-provided prompts and executed over a shared latent geological framework. This formulation allows GEM to handle multiple interpretation goals within a single architecture. As illustrated in Figure 1a, GEM supports a broad spectrum of tasks—including fault and horizon extraction, relative geological time (RGT) estimation, geobody segmentation (e.g., channels, salt, volcanic units), and property modeling (e.g., impedance, gamma ray, lithology)—all through a consistent, prompt-conditioned inference mechanism.
A defining feature of GEM is its interactive and structure-aware reasoning paradigm. Unlike conventional models that rely on static input-output mappings, GEM supports flexible prompting over input subsurface images. Experts can provide sparse inputs—such as masks, well logs, or structural sketches—and iteratively refine outputs in real time. For example, in structural interpretation or geobody segmentation, users can mark key boundaries or features on a 2D section to guide 3D inference, then adjust unsatisfactory regions through pixel-level corrections. For RGT estimation, GEM takes a vertically aligned RGT scale prompt linearly scaled from 0 to 1, and iteratively refines the prediction by injecting new prompts at regions with high misalignment (Figure 8). In property modeling, sparse well logs serve as conditioning prompts to generate geologically consistent volumes. These multimodal prompts are propagated through the latent structural framework inferred from an input subsurface image, enabling consistent and context-aware generation across tasks. Appendices Figure 13 and accompanying videos demonstrate this interactive workflow, where experts visualize, intervene, and re-infer outputs in a seamless loop. This human-in-the-loop capability allows GEM to integrate expert knowledge fluidly, improving interpretability, adaptability, and alignment with real-world geoscience practice.
@@ 2.2 Emergent Structure-Aware Reasoning in GEM
Building on this internal representation, GEM performs prompt-conditioned inference by propagating sparse user inputs—such as well logs, masks, or sketches—along the inferred geological framework. As illustrated in Figure 2 (b), one example involves using sparse well-log measurements as prompts (colored vertical trace in (b-1)) for property modeling. Because GEM can internally infer a geologic framework from the seismic image (b-1)—including features such as faults, unconformities, and marker strata—it can propagate these well-log values in a geologically consistent manner. This enables the generation of property volumes that honor both the structural context of the subsurface image and the sparse well-log conditioning. This process demonstrates how GEM leverages its internal structure awareness to produce geologically plausible outputs from minimal conditioning data.
To better understand how prompts guide internal reasoning, we freeze the pretrained GEM model and attach a lightweight multilayer perceptron (MLP) to probe its intermediate features during forward inference. As shown in Figure 2 (c), we examine how different prompt types (a 2D mask or a vertical well-log) trigger distinct patterns of feature activation. These MLP-projected visualizations reveal how prompt information propagates through successive network modules ($k_{1}$ to $k_{5}$), gradually expanding from local prompt regions into structurally coherent representations. The mask prompt activates localized features that spread laterally along a geologic stratum, while the well-log prompt initiates vertical propagation that aligns with stratigraphic continuity. This behavior mirrors expert interpretation: starting from sparse observations, GEM progressively extrapolates geologically consistent structures that respect boundaries and relationships encoded in the subsurface image. Such prompt-aware, structure-conditioned reasoning underlies GEM’s ability to unify diverse interpretation tasks—ranging from structural delineation and geobody segmentation to property modeling—within a single, generalizable framework.
@@ 2.3 Structural and Stratigraphic Interpretation
Figure 3: GEM demonstrates strong zero-shot performance across diverse subsurface interpretation tasks. (a) Quantitative evaluation of fault interpretation using intersection-over-union (IOU) on both synthetic and field datasets. GEM consistently outperforms baseline models across multiple fault types, without task-specific retraining. (b) Schematic illustrating how expert prompts (e.g., BSR and fault sketches) are combined with seismic imaging to guide GEM’s inference. (c) Zero-shot 3D interpretation of the Hikurangi subduction margin. With an average of 2 prompts per instance, GEM reconstructs a full-volume BSR surface and complex thrust fault network. (d) Interpretation of Martian polar stratigraphy from SHARAD radargrams. Panel (d-1) shows input with prompt location; panels (d-2) to (d-7) visualize surface and internal reflectors generated by GEM. The model demonstrates emergent generalization to an unseen radar modality using only a few expert-provided slice prompts. (e, f). Unconformity and RGT estimation on the Poseidon and Delft datasets. GEM produces geologically consistent RGT volumes and boundary surfaces from sparse initial constraints. More results are presented in Figure 9.
@@ 2.4 Geobody Segmentation
In the quantitative experiments on synthetic data, GEM does not exhibit a clear advantage (left side of Figure 4 (a)), similar to the case in structural interpretation tasks. Its strengths lie in generalization to field data and its powerful zero-shot learning capabilities. Particularly in the task of channel segmentation, the diversity of fluvial depositional patterns and structural complexity sharply contradicts the low information density and semantic deficiency inherent in subsurface imaging. This contradiction is a major factor limiting the generalization ability of most existing channel detection methods across different survey areas, often resulting in a typical IOU below $0.2$ for conventional AI models. Their qualitative outputs are also unreliable, frequently missing key channel structures or introducing significant false positives, as illustrated in Apendix Figure 10 (a, b).
@@ 2.5 Well-Log Prompted Geophysical Property Modeling
Figure 5: GEM enables prompt-guided, multi-property modeling with minimal well control. (a) Quantitative comparison of impedance and gamma ray modeling on synthetic (SEAM I) and field datasets (F3, Teapot). In zero-shot settings using only sparse well-log prompts, GEM outperforms specifically trained conventional models including U-Net, DeepLabv3, and HRNet. (b) Qualitative comparison on SEAM I. Panel (b-3) shows GEM’s result prompted by 4 logs; panels (b-4) to (b-6) show results from models trained with the same logs. (c) Impedance modeling (c-2) on the Teapot Dome dataset, with blind well validation shown in (c-3), where GEM better captures impedance variations with higher fidelity. (d) Multi-property modeling on the Netherlands F3, including impedance (d-2), gamma ray (d-3), and lithology (d-4), based on sparse log inputs. Panel (d-5) shows blind well validation, where GEM’s predictions remain highly consistent with ground truth. (e) Using sparse log prompts in another survey of Delft (e-1), GEM generates geologically consistent impedance (e-2) and gamma (e-3) volumes with strong agreement at non-prompted wells (e-4). More results are provided in Figure 11.
@@
GEM addresses these challenges by treating well logs as flexible prompts that introduce regional property trends, while relying on its pretrained ability to extract geological frameworks (e.g., structural boundaries and stratigraphic sequences) from the input seismic volume to guide property modeling. Instead of learning a direct mapping from data to property, GEM performs welllog-prompt-driven property completion constrained by the inferred geologic architecture.This enables the generation of property volumes that are not only geologically plausible, but also spatially consistent with seismic structures. As a result, GEM supports zero-shot modeling of diverse geophysical properties—even those not included in training—without retraining. As shown in Figure 5, we demonstrate this across four representative datasets: SEAM I (velocity), Teapot Dome (impedance), Netherlands F3 (impedance, gamma ray, lithology), and Delft (impedance and gamma ray), where GEM consistently produces high-fidelity property models from minimal well input.
To demonstrate GEM’s modeling capabilities, we compare its performance against three conventional AI models trained under supervised, multidimensional inversion frameworks. While these baselines rely on paired seismic–well data and task-specific training, GEM performs inference directly using well logs as prompts, without any retraining. In the quantitative results, for the synthetic dataset, where a complete 3D velocity volume is available, we report both semantic-level and patch-level performance metrics, including Frechet Inception Distance (FID) [59] and Structural Similarity Index (SSIM) [60]. For the field datasets, that lack full-volume ground truth, evaluation is conducted using pixel-level Mean Absolute Error (MAE) calculated at blind well locations—that is, wells excluded from both the training set of conventional models and the prompt input to GEM. This ensures a fair and consistent comparison of generalization performance across methods. The quantitative metrics in Figure 5a clearly demonstrate GEM’s superior performance in property modeling: it achieves consistently higher SSIM and lower MAE on synthetic data, and lower MAE on field datasets, outperforming conventional supervised models across all test cases.

### v3

# Geological Everything Model 3D: A Physics-informed Promptable Foundation Model for Unified and Zero-Shot Subsurface Understanding

@@ Abstract
Understanding Earth’s subsurface is critical for energy transition, natural hazard mitigation, and planetary science. Yet subsurface analysis remains fragmented, with separate models required for structural interpretation, stratigraphic analysis, geobody segmentation, and physical property modeling—each tightly coupled to specific data distributions and task formulations. We introduce the Geological Everything Model 3D (GEM), a physics-informed, promptable foundation model that reformulates these tasks as conditional inference along latent structural frameworks derived from subsurface imaging. By embedding geological and physical priors into its generative architecture, GEM integrates domain knowledge with machine learning to ensure structurally consistent and geologically plausible outputs. This formulation moves beyond task-specific models by enabling a shared inference mechanism, where GEM propagates human-provided prompts—such as well logs, masks, or structural sketches—through inferred structural frameworks to produce coherent subsurface interpretations. Through this mechanism, GEM achieves zero-shot generalization across tasks and modalities with heterogeneous prompt types, without retraining for new tasks or data sources. This capability emerges from a two-stage training process that combines large-scale self-supervised representation learning on over 500 field seismic volumes with adversarial fine-tuning using mixed prompts and labels across structural, stratigraphic, geobody, and property domains. GEM demonstrates broad applicability across surveys and tasks, including Martian radar stratigraphy analysis, structural interpretation in subduction zones, seismic stratigraphic interpretation, geobody delineation, and physical property modeling. By bridging expert knowledge with physics-informed generative reasoning in a structurally aware manner, GEM lays the foundation for scalable, human-in-the-loop geophysical AI—transitioning from fragmented workflows to a vertically integrated, promptable reasoning system.
@@ 1 Introduction
Subsurface understanding is foundational to some of society’s most pressing challenges: sustainable energy exploration [1, 2], geological carbon storage [3, 4], natural hazard prediction [5, 6] and stratigraphy [7, 8]. In recent years, Artificial Intelligence (AI) has emerged as a transformative paradigm for subsurface analysis [9, 10, 11, 12, 13]. However, current implementations remain fragmented, brittle, and narrowly scoped. Most workflows depend on task-specific models—each trained on narrow datasets and bound to a single survey region—to perform structural interpretation [14, 15, 16], stratigraphic analysis[17, 18, 19], geobody segmentation [20, 21, 22, 23] or physical property modeling [24, 25, 26, 27]. This fragmentation hinders generalization and requires costly retraining for every new task or area. While foundation models have revolutionized vision and language[28, 29, 30, 31, 32], no such model has emerged for subsurface imaging. This is not simply a matter of scale, but of epistemology: geological imaging data is sparse, indirect, and structurally entangled—lacking the high-density semantic cues that foundation models in other fields rely on. Bridging this domain gap calls for rethinking how we represent, prompt, and reason about the planetary interior—not as a patchwork of disconnected labels, but as a unified structure-conditioned generative process.
This study situates GEM within physics informed deep learning by encoding geological and physical priors such as stratigraphic ordering, structural continuity and physical property logging as differentiable, structure aware constraints in both the generative architecture and the learning objectives. As a result, the model delivers reliable predictive modeling under sparse supervision and noisy, indirect observations and enables expert guided, data driven scientific discovery through prompt conditioned inference over coherent latent structural representations.
@@
In response to these limitations, we introduce the Geological Everything Model 3D (GEM), a unified generative framework that reformulates subsurface interpretation as prompt-conditioned inference over latent structural representations. Instead of treating tasks like structural interpretation, geobody segmentation, or physical property modeling as separate problems requiring bespoke models, GEM abstracts them as different forms of dimensional completion guided by human-provided prompts—such as well logs, masks, or structural sketches. At its core, GEM learns to propagate these prompts through a structurally coherent latent space derived from the input subsurface volume, enabling geologically consistent outputs across diverse task types. This formulation unifies a wide range of interpretation goals within a single model, while maintaining flexibility for expert interaction and scalability across surveys. By shifting the modeling paradigm from task-specific classification to structure-aware generation, GEM offers a foundation for zero-shot, prompt-driven, human-in-the-loop geological reasoning.
@@
GEM is trained in two stages to enable broad generalization without task-specific retraining. First, a self-supervised pretraining stage leverages masked representation learning over more than 500 field seismic volumes to learn structure-sensitive features without requiring labels. Then, adversarial fine-tuning is performed using synthetic data with diverse prompt-label pairs across structural, stratigraphic, geobody, and physical property domains. This combination allows GEM to align prompts with latent geological frameworks, supporting flexible input types—including lines, masks, and well logs—and producing task-consistent outputs with minimal supervision. Once trained, GEM generalizes across surveys, data modalities, and interpretation goals, achieving zero-shot performance without architectural modification or additional fine-tuning. We evaluate GEM across a diverse set of tasks and geological settings, demonstrating broad applicability and strong zero-shot generalization. Without retraining, GEM performs high-resolution Martian radar stratigraphy analysis, interprets complex 3D fault systems in subduction zones, conducts full seismic stratigraphic interpretation, segments geobodies such as salt domes and channels, and models properties across surveys with sparse well control. These applications span planetary exploration, structural geology and physical property modeling—tasks that traditionally require disconnected models, domain-specific workflows, and extensive retraining at each stage. In contrast, GEM consolidates structural interpretation, stratigraphic analysis, geological modeling, and property prediction within a single generative framework. This unified approach eliminates the need for task-specific pipelines and mitigates error accumulation across stages, offering a more stable, scalable, and interpretable alternative. By integrating expert prompts with structure-aware reasoning, GEM bridges geoscientific insight and generative modeling, establishing a new paradigm for human-guided subsurface AI. This work makes three key contributions:
@@
  • First, it proposes a unified generative mechanism for subsurface interpretation by reframing structural delineation, full stratigraphic analysis, geobody segmentation, and physical property modeling as prompt-guided completions over latent geological frameworks derived from imaging data. This formulation enables structure-aware reasoning across tasks within a shared representation space.
@@
  • Third, it demonstrates GEM ’s strong zero-shot generalization and broad applicability across surveys, modalities, and interpretation goals, including Martian radar stratigraphy, 3D fault interpretation in subduction zones, full seismic stratigraphic interpretation, geobody segmentation, and physical property modeling with sparse well control—often superior to task-specific supervised models.
@@ 2.1 GEM: A Unified and Prompt-Driven Reasoning Framework for Subsurface imaging
GEM is designed as a unified, prompt-driven foundation model for diverse subsurface imaging tasks. It reformulates structural interpretation, stratigraphic analysis, geobody segmentation, and physical property modeling as conditional generative processes, driven by sparse user-provided prompts and executed over a shared latent geological framework. This formulation allows GEM to handle multiple interpretation goals within a single architecture. As illustrated in Figure 1a, GEM supports a broad spectrum of tasks—including fault and horizon extraction, relative geological time (RGT) estimation, geobody segmentation (e.g., channels, salt, volcanic units), and physical property modeling (e.g., impedance, gamma ray, lithology)—all through a consistent, prompt-conditioned inference mechanism.
A defining feature of GEM is its interactive and structure-aware reasoning paradigm. Unlike conventional models that rely on static input-output mappings, GEM supports flexible prompting over input subsurface images. Experts can provide sparse inputs—such as masks, well logs, or structural sketches—and iteratively refine outputs in real time. For example, in structural interpretation or geobody segmentation, users can mark key boundaries or features on a 2D section to guide 3D inference, then adjust unsatisfactory regions through pixel-level corrections. For RGT estimation, GEM takes a vertically aligned RGT scale prompt linearly scaled from 0 to 1, and iteratively refines the prediction by injecting new prompts at regions with high misalignment (Figure 8). In physical property modeling, sparse well logs serve as conditioning prompts to generate geologically consistent volumes. These multimodal prompts are propagated through the latent structural framework inferred from an input subsurface image, enabling consistent and context-aware generation across tasks. Appendices Figure 13 and accompanying videos demonstrate this interactive workflow, where experts visualize, intervene, and re-infer outputs in a seamless loop. This human-in-the-loop capability allows GEM to integrate expert knowledge fluidly, improving interpretability, adaptability, and alignment with real-world geoscience practice.
@@ 2.2 Emergent Structure-Aware Reasoning in GEM
Building on this internal representation, GEM performs prompt-conditioned inference by propagating sparse user inputs—such as well logs, masks, or sketches—along the inferred geological framework. As illustrated in Figure 2 (b), one example involves using sparse well-log measurements as prompts (colored vertical trace in (b-1)) for physical property modeling. Because GEM can internally infer a geologic framework from the seismic image (b-1)—including features such as faults, unconformities, and marker strata—it can propagate these well-log values in a geologically consistent manner. This enables the generation of property volumes that honor both the structural context of the subsurface image and the sparse well-log conditioning. This process demonstrates how GEM leverages its internal structure awareness to produce geologically plausible outputs from minimal conditioning data.
To better understand how prompts guide internal reasoning, we freeze the pretrained GEM model and attach a lightweight multilayer perceptron (MLP) to probe its intermediate features during forward inference. As shown in Figure 2 (c), we examine how different prompt types (a 2D mask or a vertical well-log) trigger distinct patterns of feature activation. These MLP-projected visualizations reveal how prompt information propagates through successive network modules ($k_{1}$ to $k_{5}$), gradually expanding from local prompt regions into structurally coherent representations. The mask prompt activates localized features that spread laterally along a geologic stratum, while the well-log prompt initiates vertical propagation that aligns with stratigraphic continuity. This behavior mirrors expert interpretation: starting from sparse observations, GEM progressively extrapolates geologically consistent structures that respect boundaries and relationships encoded in the subsurface image. Such prompt-aware, structure-conditioned reasoning underlies GEM’s ability to unify diverse interpretation tasks—ranging from structural delineation and geobody segmentation to physical property modeling—within a single, generalizable framework.
@@ 2.3 Structural and Stratigraphic Interpretation
Figure 3: GEM demonstrates strong zero-shot performance across diverse subsurface interpretation tasks. (a) Quantitative evaluation of fault interpretation using intersection-over-union (IOU) on both synthetic and field datasets. GEM consistently outperforms baseline models across multiple fault types, without task-specific retraining. (b) Schematic illustrating how expert prompts (e.g., BSR and fault sketches) are combined with seismic imaging to guide GEM’s inference. (c) Zero-shot 3D interpretation of the Hikurangi subduction margin. With an average of 2 prompts per instance, GEM reconstructs a full-volume BSR surface and complex thrust fault network. (d) Interpretation of Martian polar stratigraphy from SHARAD radargrams. Panel (d-1) shows input with prompt location; panels (d-2) to (d-7) visualize surface and internal reflectors generated by GEM. The model demonstrates emergent generalization to an unseen radar modality using only a few expert-provided mask prompts. (e, f). Unconformity and RGT estimation on the Poseidon and Delft datasets. GEM produces geologically consistent RGT volumes and boundary surfaces from sparse initial constraints. More results are presented in Figure 9.
@@ 2.4 Geobody Segmentation
In the quantitative experiments on synthetic data, GEM does not exhibit a clear advantage (left side of Figure 4 (a)), similar to the case in structural interpretation tasks. Its strengths lie in generalization to field data and its powerful zero-shot learning capabilities. Particularly in the task of channel segmentation, the diversity of fluvial depositional patterns and structural complexity sharply contradicts the low information density and semantic deficiency inherent in subsurface imaging. This contradiction is a major factor limiting the generalization ability of most existing channel detection methods across different survey areas, often resulting in a typical IOU below $0.2$ for conventional AI models. Their qualitative outputs are also unreliable, frequently missing key channel structures or introducing significant false positives, as illustrated in Appendices Figure 10 (a, b).
@@ 2.5 Well-Log Prompted Geophysical Property Modeling
Figure 5: GEM enables prompt-guided, multi-physical-property modeling with minimal well control. (a) Quantitative comparison of impedance and gamma ray modeling on synthetic (SEAM I) and field datasets (F3, Teapot). In zero-shot settings using only sparse well-log prompts, GEM outperforms specifically trained conventional models including U-Net, DeepLabv3, and HRNet. (b) Qualitative comparison on SEAM I. Panel (b-3) shows GEM’s result prompted by 4 logs; panels (b-4) to (b-6) show results from models trained with the same logs. (c) Impedance modeling (c-2) on the Teapot Dome dataset, with blind well validation shown in (c-3), where GEM better captures impedance variations with higher fidelity. (d) Multi-physical-property modeling on the Netherlands F3, including impedance (d-2), gamma ray (d-3), and lithology (d-4), based on sparse log inputs. Panel (d-5) shows blind well validation, where GEM’s predictions remain highly consistent with ground truth. (e) Using sparse log prompts in another survey of Delft (e-1), GEM generates geologically consistent impedance (e-2) and gamma (e-3) volumes with strong agreement at non-prompted wells (e-4). More results are provided in Figure 11.
@@
GEM addresses these challenges by treating well logs as flexible prompts that introduce regional property trends, while relying on its pretrained ability to extract geological frameworks (e.g., structural boundaries and stratigraphic sequences) from the input seismic volume to guide physical property modeling. Instead of learning a direct mapping from data to property, GEM performs welllog-prompt-driven property completion constrained by the inferred geologic architecture.This enables the generation of property volumes that are not only geologically plausible, but also spatially consistent with seismic structures. As a result, GEM supports zero-shot modeling of diverse geophysical properties—even those not included in training—without retraining. As shown in Figure 5, we demonstrate this across four representative datasets: SEAM I (velocity), Teapot Dome (impedance), Netherlands F3 (impedance, gamma ray, lithology), and Delft (impedance and gamma ray), where GEM consistently produces high-fidelity property models from minimal well input.
To demonstrate GEM’s modeling capabilities, we compare its performance against three conventional AI models trained under supervised, multidimensional inversion frameworks. While these baselines rely on paired seismic–well data and task-specific training, GEM performs inference directly using well logs as prompts, without any retraining. In the quantitative results, for the synthetic dataset, where a complete 3D velocity volume is available, we report both semantic-level and patch-level performance metrics, including Frechet Inception Distance (FID) [59] and Structural Similarity Index (SSIM) [60]. For the field datasets, that lack full-volume ground truth, evaluation is conducted using pixel-level Mean Absolute Error (MAE) calculated at blind well locations—that is, wells excluded from both the training set of conventional models and the prompt input to GEM. This ensures a fair and consistent comparison of generalization performance across methods. The quantitative metrics in Figure 5a clearly demonstrate GEM’s superior performance in physical property modeling: it achieves consistently higher SSIM and lower MAE on synthetic data, and lower MAE on field datasets, outperforming conventional supervised models across all test cases.